OpenStudy (anonymous):

Find the probability for the experiment of drawing two marbles (without replacement) from a bag containing four green, five yellow, and two red marbles such that both marbles are yellow.

OpenStudy (anonymous):

\[\frac{5}{11}\times \frac{4}{10}\]